how would you solve
y=-x+3
y=x+1

Answer by suruman(21)

y = -x + 3 -----eqn(1)
y = x + 1 -----eqn(2)
Adding the above equations,
y + y = -x + 3 + x + 1
2*y = 4
=> y = 2
Substituting value of y in eqn(1),
2 = -x + 3
x = 3 - 2 = 1
Thus the solution pair for the set of equations eqn(1) and eqn(2) is :
(x,y) = (1,2).